# Black-box optimization and quantum annealing for filtering out mislabeled training instances

### v1

@@ Results
Figure 7: Characteristics of the solutions found by each sampler in 32 runs. (a) The loss value of the best solution found by each sampler in each run is represented as a single point. Its distribution over 32 runs is shown as a box plot for each sampler. The whiskers extend to the farthest points within 1.5 times the inter-quartile range from the nearest hinge of the box plot. (b) Hamming distance between the theoretical solution and the best solution found by each sampler in each run is shown in the same format.
Fig. 6 also shows the interesting difference in how 512 samples are utilized by each sampler in each optimization step. Firstly, in case of OpenJij SQA sampler, the optimal sample with lowest energy in 512 samples is always accepted in each optimization step. This is due to the fact that both the variance and the mean of energy distribution of samples obtained by OpenJij SQA sampler is always higher than ones obtained by Neal SA sampler or D-Wave QA clique sampler as shown in Fig. 8. Secondly, in case of Neal SA sampler, most of 512 samples obtained in each step are the same and have equal energy as shown in Fig. 8(b) (In this particular run, they are all equal). Therefore, there is a higher chance for the optimal sample in a certain step to be already seen before, and it cause the higher chance of random samples to be selected (See the higher probability of random samples marked by black cross shown in Fig. 6(b). Thirdly, in the case of D-Wave QA clique sampler, suboptimal sample with sufficiently low energy is fully utilized to find lower energy solution in each optimization step as shown in Fig. 6(c).
@@
Figure 8: Energy distribution of 512 samples obtained in each optimization step of three different samplers: (a) OpenJij SQA sampler, (b) Neal SA sampler, and (c) D-Wave QA clique sampler. The top figure for each sampler shows the raw energy values of 512 samples obtained in each step over the entire 256 steps in the optimization phase. The mean energy value of 512 samples acquired in each step is shown with the blue line. The bottom figure for each sampler shows the standard deviation of the energy values of 512 samples acquired in each step. Please note the energy scales of Neal SA sampler and and D-Wave QA clique sampler are the same but different from the one used for the OpenJij SQA sampler.
@@
In addition to achieving high solution quality, the D-Wave QA clique sampler demonstrated significantly faster sampling time compared to other samplers when using the default parameter settings. Sampling time of D-Wave QA clique sampler is about 10 times faster than one of Neal SA sampler and is about 100 times faster than one of OpenJij SQA sampler as shown in Fig. 9.
@@
Figure 9: Sampling time statistics of three samplers. (a) Sampling times in each step are compared for the three samplers. The 95% confidence intervals obtained after running 32 runs are shown with transparent bands around the mean line plots. (b) The standard error of the mean over 32 runs, where the mean is taken over 256 steps of sampler’s sampling times. The whiskers extend to the farthest points within 1.5 times the inter-quartile range from the nearest hinge of the box plot.
@@
The proposed algorithm successfully identified mislabeled instances contained in the contaminated training dataset as shown in the previous section. Because our task is designed to be impossible to solve above the chance level without proper identification of mislabeled instances, the successful improvement in the generalization performance of the trained model, which is reflected in the test error, indicates that the proposed algorithm is proved to be effective in identifying mislabeled instances. Our approach to the problem of removing mislabeled instances became possible by combining the surrogate model-based BBO and quantum annealer with clearly defined loss function using noise-free validation dataset.
We have obtained several important insights from the results. First, we have observed that the proposed algorithm removed mislabeled instances that have more detrimental effect on the model’s generalization performance with higher probability. This favorable tendency likely arises from leveraging the validation error–directly tied to generalization performance–as the objective function in the surrogate model-based BBO. In our task, mislabeled instances with lower input entropy tend to exert a stronger negative influence on generalization performance and are therefore more likely to be removed. This suggests that the proposed algorithm inherently possesses the capability to prioritize the removal of mislabeled instances that exert a more detrimental impact on the generalization ability of the trained model. Therefore, the proposed algorithm is also likely be effective for removing mislabeled instances with complex input features such as images or texts. Second, we have observed that the proposed algorithm effectively leveraged the sample diversity provided by the quantum annealer, driven in part by quantum fluctuations, to achieve a lower validation loss. Quantum annealers are capable of generating diverse solutions rapidly due to their inherent quantum fluctuations, which facilitates exploration of the solution space. As a result, the combination of the quantum annealer and the surrogate model-based BBO may prove valuable for addressing combinatorial optimization problems with a vast solution space, extending beyond noise removal to other problem setups.
The proposed algorithm has multiple future directions to explore. Here in this study, we have focused on the problem of discarding unimportant instances from a given dataset. However, we can easily extend the proposed algorithm to simultaneously discard unimportant features as well as instances by allocating binary nodes over features as well as instances. It is of course possible to use the proposed method just for feature selection by allocating binary nodes only over features. The objective function remains the same, and we can automatically find relevant features for building the model with high generalization performance. This is the first interesting future direction to explore. Another interesting future direction is to evaluate the proposed algorithm on real-world datasets. In this study, we have only evaluated the proposed algorithm on synthetic datasets to show the effectiveness of the proposed algorithm for removing mislabeled dataset in the hardest case. However, it is important to evaluate the proposed algorithm on real-world datasets to investigate the effectiveness of the proposed algorithm in practical applications. The real world datasets contains different degrees of noises in labels and attributes, and it is important to evaluate how effective the proposed algorithm on these real-world datasets. Also, we used a linear regression model as a base model in this study, but any kind of machine learning models can be used as a base model. Due to the iterative nature of the proposed algorithm, not just sampling time but also the training time of the base model becomes important issue when handing real-world datasets. Therefore, it is important to see how much performance improvement in terms of generalization can be achieved by the proposed algorithm, given the fixed real-world datasets at hand. Yet another interesting future direction is to apply the proposed algorithm to unsupervised learning problems. In this study, we have mainly focused on the supervised learning problem aimed at removing mislabeled instances from a given dataset. However, the proposed algorithm can be also applied to unsupervised learning settings by estimating the suitable objective function such as negative log-partition function of unlabeled observations by the surrogate model. The proposed algorithm is likely to be used to remove unimportant instances or features in the dataset in unsupervised learning settings.
Although the proposed algorithm has shown promising results in a well-controlled setting, there are several limitations in this study. First limitation comes from the fact that the task is artificially created and relatively small in its size. The task is designed to be simple enough to be solved by the logistic regression model, and the number of data points is 128, which is not that large in usual machine learning setting where the simple MNIST dataset has 60,000 training data points. Second limitation comes from the fact that we used the logistic regression model with $L_{2}$ regularization as a base model. This choice comes from our prior knowledge for the noisy majority bit task, but it is possible that the LASSO model, which can automatically select important features, might be more suitable for many real-world applications. Third limitation comes from the fact that we have not evaluated the proposed algorithm on large-scale datasets. The proposed algorithm has shown promising results in the noisy majority bit task with 128 instances, but it is possible that the proposed algorithm might not work well on large-scale datasets. Finally, the proposed algorithm is limited by the number of qubits in the quantum annealer. So, if the number of data points exceeds the number of qubits in the quantum annealer, the problem cannot be embedded into the quantum annealer, and the proposed algorithm cannot be applied to the problem.
@@
In this study, we have proposed a novel algorithm capable of removing mislabeled instances from a given training dataset. This algorithm estimates the loss against a correctly-labeled validation dataset using surrogate model-based BBO followed by postprocessing. Additionally, we have demonstrated that the proposed algorithm can be further enhanced in both solution quality and sampling speed by leveraging the unique characteristics of D-Wave Systems’ quantum annealer, such as its sampling diversity and rapid sampling capability. The proposed algorithm provides a new direction for improving the quality of training datasets by combining the power of BBO and quantum annealing. This combination offers a promising approach to tackling the long-standing challenge of data quality improvement. In the future, it is important to evaluate the proposed algorithm on real-world datasets to assess its effectiveness in practical applications. This evaluation will help determine the tangible benefits and potential limitations of the algorithm when applied to complex and diverse data scenarios.
@@ Acknowledgements
We would like to extend our sincere gratitude to all participants of Quantum Infinity for You, The First Session with You, in Kumamoto (QI4U). In particular, we would like to thank Tadashi Kadowaki for guiding our research in the initial phase and Akira Sasao for organizing the event.

### v2

@@ Results
Figure 7: Characteristics of the solutions found by each sampler in 32 runs. (a) The surrogate loss value, which is the log-transformed validation loss, of the best solution found by each sampler in each run is represented as a single point. Its distribution over 32 runs is shown as a box plot for each sampler. The whiskers extend to the farthest points within 1.5 times the inter-quartile range from the nearest hinge of the box plot. (b) Hamming distance between the theoretical solution and the best solution found by each sampler in each run is shown in the same format. This Hamming distance is equivalent to the prediction accuracy of the BBO model when multiplied by the size of the training dataset. After collecting 96 solutions obtained by 3 different samplers over 32 runs, the D-Wave QA sampler predominantly produced the top-performing solutions in terms of both log-loss error and Hamming distance.
Fig. 6 also shows the interesting difference in how 512 samples are utilized by each sampler in each optimization step. Firstly, in case of OpenJij SQA sampler, the optimal sample with lowest energy in 512 samples is always accepted in each optimization step. This is due to the fact that both the variance and the mean of energy distribution of samples obtained by OpenJij SQA sampler is always higher than ones obtained by Neal SA sampler or D-Wave QA clique sampler as shown in Fig. 8. Secondly, in case of Neal SA sampler, most of 512 samples obtained in each step are the same and have equal energy as shown in Fig. 8(b) (In this particular run, they are all equal). Therefore, there is a higher chance for the optimal sample in a certain step to be already seen before, and it causes a higher chance of random samples to be selected (See the higher probability of random samples marked by black cross shown in Fig. 6(b). Thirdly, in the case of D-Wave QA clique sampler, suboptimal sample with sufficiently low energy is fully utilized to find lower energy solution in each optimization step as shown in Fig. 6(c).
@@
Figure 8: Energy distribution of 512 samples obtained in each optimization step of three different samplers: (a) the OpenJij SQA sampler, (b) the Neal SA sampler, and (c) the D-Wave QA clique sampler. The top figure for each sampler shows the raw energy values of 512 samples obtained in each step over the entire 256 steps in the optimization phase. The mean energy value of 512 samples acquired in each step is shown with the blue line. The bottom figure for each sampler shows the standard deviation of the energy values of 512 samples acquired in each step. Please note the energy scales of the Neal SA sampler and the D-Wave QA clique sampler are the same but different from the one used for the OpenJij SQA sampler. As illustrated in these figures, the three samplers have distinct sampling behaviors: (a) the OpenJij SQA sampler produces highly variable samples with elevated energy, (b) the Neal SA sampler generates low-energy but identical samples, (c) the D-Wave QA clique sampler maintains a moderate level of sample variability even in the low-energy regime. This distinctive characteristic of the D-Wave’s physical sampler suggests that it may promote exploration in the vicinity of the global minimum.
@@
In addition to achieving high solution quality, the D-Wave QA clique sampler demonstrated significantly faster sampling time compared to other samplers when using the default parameter settings. Sampling time of D-Wave QA clique sampler is about 10 times faster than that of the Neal SA sampler and about 100 times faster than that of the OpenJij SQA sampler, as shown in Fig. 9. To assess the robustness of these differences, we conducted pairwise Mann–Whitney–Wilcoxon tests across the 32 runs. All pairwise comparisons yielded highly significant results ($p=6.511\times 10^{-12}$, $U=1.024\times 10^{3}$), confirming that the observed superiority of the D-Wave QA clique sampler over classical samplers is statistically significant.
@@
Figure 9: Sampling time statistics of three samplers. (a) Sampling times in each step are compared for the three samplers. The reported time is measured in wall-clock time and includes the time required for embedding. The 95% confidence intervals obtained after running 32 runs are shown with transparent bands around the mean line plots. (b) The standard error of the mean over 32 runs, where the mean is taken over 256 steps of sampling times of each sampler. The whiskers extend to the farthest points within 1.5 times the inter-quartile range from the nearest hinge of the box plot. Overall, the mean sampling times of the OpenJij SQA sampler, the Neal SA sampler, and the D-Wave QA clique sampler are 1.196, 0.104, and 0.011 seconds, respectively. These results demonstrate that the D-Wave QA clique sampler achieves sampling times more than an order of magnitude faster than the other two classical samplers, highlighting its potential to obtain higher-quality solutions in substantially less time.
@@
The proposed algorithm successfully identified mislabeled instances contained in the contaminated training dataset, as demonstrated by the substantial improvements in generalization performance. Since the noisy majority-bit task is unsolvable above chance level without removing mislabeled instances, the improved test performance clearly indicates successful noise removal. This success was made possible by directly optimizing validation loss via surrogate model-based black-box optimization (BBO) with postprocessing and leveraging the unique sampling capabilities of quantum annealing.
One of the key insights from this study is the observed prioritization of harmful mislabeled instances. The algorithm was more likely to remove mislabeled samples with low input entropy, which tend to exert a stronger negative effect on generalization. This behavior naturally emerged from the use of validation loss as the optimization objective and reflects the ability of the surrogate model-based BBO to exploit structural patterns in the data that correlate with label correctness.
More critically, we found that quantum annealing offers distinct advantages in this specific noise-removal task, beyond what is achievable with classical samplers. In high-dimensional combinatorial problems such as instance selection (where the search space contains $2^{n}$ subsets), exploration efficiency is paramount. Classical samplers like simulated annealing (SA) and simulated quantum annealing (SQA) often converge to narrow solution basins and suffer from sample degeneracy, limiting the breadth of search.
In contrast, D-Wave’s quantum annealer consistently produced diverse and energetically favorable solutions in each optimization step (Fig. 8(c)), resulting in lower validation loss (Fig. 6(c)) and better alignment with the theoretical optimum (Fig. 7). This is likely due to quantum tunneling effects, which enable the system to escape shallow local minima and explore disconnected regions of the solution space more effectively than classical thermal processes.
Moreover, the reduced reliance on random sampling in later stages of optimization (Fig. 6(c)) and the high likelihood to find the lower surrogate losses (Fig. 7(a)) highlight the practical advantage of quantum annealing in iterative optimization under sampling constraints. These benefits were achieved with significantly lower sampling time compared to classical samplers (Fig. 9), enhancing both solution quality and computational efficiency.
Taken together, these results suggest that quantum annealing is particularly well-suited for subset selection tasks in machine learning, especially when the solution space is vast and multi-modal. Its ability to produce diverse, high-quality candidates at low latency offers a compelling alternative to traditional methods that rely on slower or more redundant sampling strategies.
Looking ahead, the proposed algorithm can be extended in several directions. First, by introducing binary selection over features as well as instances, it can support joint instance and feature selection within the same framework. Second, while this study focused on synthetic data for clarity, evaluating the approach on real-world datasets with naturally occurring label noise is essential for assessing practical utility. Finally, the method may be adapted to unsupervised learning tasks by designing suitable surrogate objectives, such as approximations to negative log-likelihood.
While the proposed algorithm demonstrates promising results in a well-controlled setting, there are several limitations in this study. First, the experimental task–namely the noisy majority-bit task–is synthetic and relatively small in scale, consisting of only 128 instances. This simplified setting was deliberately chosen to allow clear analysis and control over the noise structure. However, we recognize that such a task may not fully reflect the complexity of real-world datasets, such as CIFAR-10N/100N [38] or Clothing1M [39], where label noise occurs in more diverse and less structured forms. Evaluation on large-scale, real-world benchmarks is necessary to assess its practical utility in future studies. Second, we adopted a logistic regression model with $L_{2}$ regularization as a base learner, given prior knowledge about the task. While this choice was sufficient for our controlled study, alternative models such as LASSO or deep neural networks may be more suitable in practical real-world applications. Finally, the applicability of the quantum annealer is constrained by its hardware limitations, particularly the number of available qubits. If the size of the training dataset exceeds the embedding capacity of the annealer, alternative sampling strategies or a hybrid approach must be considered [22]. In addition, current quantum annealers suffer from sparse connectivity between qubits, which often requires complex minor embedding procedures that introduce additional overhead and reduce effective problem size [40]. This embedding overhead, combined with limited connectivity, can significantly degrade solution quality and sampling efficiency. Moreover, thermal noise and analog control errors in the annealing process can lead to deviations from the ideal ground-state solution, further limiting reliability in practical use cases [41]. Despite these limitations, we believe the proposed study provides a useful proof-of-concept that motivates further investigation into robust sample selection in noisy-label settings, especially in combinatorial spaces where standard gradient-based approaches may be less effective.
@@
In this study, we have proposed a novel algorithm capable of removing mislabeled instances from a given training dataset. This algorithm estimates the loss against a correctly-labeled validation dataset using surrogate model-based BBO followed by postprocessing. Additionally, we have demonstrated that the proposed algorithm can be further enhanced in both solution quality and sampling speed by leveraging the unique characteristics of D-Wave Systems’ quantum annealer, such as its sampling diversity and rapid sampling capability. The proposed algorithm provides a new direction for improving the quality of training datasets by combining the power of postprocessing-equipped BBO and quantum annealing. This combination offers a promising approach to tackling the long-standing challenge of data quality improvement. In the future, it is important to evaluate the proposed algorithm on real-world datasets to assess its effectiveness in practical applications. This evaluation will help determine the tangible benefits and potential limitations of the algorithm when applied to complex and diverse data scenarios.
Data availability
The datasets used and/or analyzed during the current study are available from the corresponding author on reasonable request.
@@ Acknowledgements
We gratefully acknowledge the support provided by the SIP/BRIDGE program. Additionally, we extend our sincere gratitude to all participants of Quantum Infinity for You: The First Session with You, held in Kumamoto (QI4U). In particular, we would like to express our deep appreciation to Dr. Tadashi Kadowaki for his invaluable guidance during the initial phase of our research and to Dr. Akira Sasao for his dedicated efforts as the local organizer of the event.
Funding
This work was financially supported by the programs for bridging the gap between R&D and IDeal society (Society 5.0) and Generating Economic and social value (BRIDGE) and Cross-ministerial Strategic Innovation Promotion Program (SIP) from the Cabinet Office (23836436).